OpenStudy (anonymous):

Amanda drove 225 miles using 8 gallons of gas. At this rate, how many gallons of gas would she need to drive 315 miles?

OpenStudy (ash2326):

\[\text{225 miles} \longrightarrow \text{8 gallons for gas}\] \[\text{For 1 mile} \longrightarrow \frac{8}{225}\text{ gallons for gas}\] \[\text{For 315 miles} \longrightarrow \frac{8}{225}\times 315\text{ gallons for gas}\]

OpenStudy (anonymous):

\(\ \huge \frac{225mi}{8gallons} = \frac{315mi}{x gallons} \). Solving for x, \(\ \huge 225x=2520 \). Therefore she will need \(\ \huge \frac{2520}{225} = \frac{56}{5} gallons, or 11.2 gallons \).